OpenStudy (anonymous):

Solve for x. logx+ log(x-1) = log(4x)

OpenStudy (ikram002p):

now convert the lft hand side to one log logx+ log(x-1) = log(4x) log(x*(x-1))=log(4x) since the base of both log are the same x(x-1)=4x x^2-x=4x x^2-5x=0 x(x-5)=0 now solve for x x=0 x-5=0.......x=5 got it or still cofused ?!

OpenStudy (anonymous):

got it :) thank you!!!